Find an equation of the line passing through
the point (7, 6) with slope m = 2.

Answer by stanbon(75887)

Find an equation of the line passing through
the point (7, 6) with slope m = 2.
------
Form: y = mx + b
6 = 2*7 + b
b = 6-14
b = -8
-------
Equation:
y = 2x-8
==========
Cheers,
Stan H.